OpenStudy (anonymous):

PLEASE HELP WILL GIVE A MEDAL Use graphs and tables to find the limit and identify any vertical asymptotes of

OpenStudy (solomonzelman):

you want the numerical method, I guess?

OpenStudy (solomonzelman):

numerical *approach.... k, shoot the question.

OpenStudy (solomonzelman):

So you want to fill in the table. \(\large\color{black}{ \huge{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr f(x) \\ \hline \scr~~~8.0~ & \scr 0 \\ \hline \scr~~~8.5~ & \scr 0 \\ \hline \scr~~~8.9~ & \scr 0 \\ \hline \scr~~~8.99~ & \scr 0 \\ \hline \scr~~~8.999~ & \scr 0 \\ \hline \end{array} } }\)

OpenStudy (anonymous):

i don't understand @SolomonZelman

OpenStudy (solomonzelman):

that would be the table to fill in to find \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~9^-}f(x)}\) .

OpenStudy (solomonzelman):

You are basically finding the output, for numbers that approach 9 from the left side.

OpenStudy (solomonzelman):

it doesn't have to use those specific numbers, but you would want to choose a bunch of numbers that are close to 9, and getting closer and closer to 9.

OpenStudy (anonymous):

but dont you just keep getting numbers close to -1 for f(x)

OpenStudy (solomonzelman):

you want the numbers close to 9, because you are finding the \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~9^-}~\frac{1}{x-9}}\)

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\frac{1}{\color{blue}{8}-9}=\frac{1}{-1}=-1}\) \(\large\color{slate}{\displaystyle\frac{1}{\color{blue}{8.5}-9}=\frac{1}{-0.5}=-2}\) \(\large\color{slate}{\displaystyle\frac{1}{\color{blue}{8.9}-9}=\frac{1}{-0.1}=-10}\) \(\large\color{slate}{\displaystyle\frac{1}{\color{blue}{8.99}-9}=\frac{1}{-0.01}=-100}\) \(\large\color{slate}{\displaystyle\frac{1}{\color{blue}{8.999}-9}=\frac{1}{-0.001}=-1000}\)

OpenStudy (solomonzelman):

See, as I am plugging numbers closer and closer to 9 (from the left side) we get an input that closer and closer approaches negative infinity.

OpenStudy (anonymous):

wait so 9 or -9

OpenStudy (solomonzelman):

you are finding \(\large\color{slate}{\displaystyle\lim_{x \rightarrow ~9^-}\frac{1}{x-9}}\)

OpenStudy (solomonzelman):

I was plugging numbers for x (that approach 9 from the left) , that is all.

OpenStudy (anonymous):

so the vertical asymptote is 9?

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

oh okay thank you!

OpenStudy (solomonzelman):

do you fully understand why though?

OpenStudy (anonymous):

yes through plugging the numbers in we see that

OpenStudy (solomonzelman):

basically, the idea here is: \(\large\color{slate}{\displaystyle\frac{1}{x-9}}\). as x is approaching 9 from the right, the x will be a little greater than 9. NOW, the closer x gets to 9 (from right side) thesmaller decimal you get on the denominator.

OpenStudy (solomonzelman):

1 / infinitely small (positive) decimal is same thing as infinity

OpenStudy (solomonzelman):

So your graph, you can already make.